Astronomy Report – March 13th

Crew 223 Astronomy Report Mar 13 2020

Name: Florian Delpech

Crew: 223

Date: 3/13/2020


Robotic Telescope Requested (choose one) MDRS-14

Objects to be Imaged this Evening: N/A because of the bad weather

Images submitted with this report: 1 picture of the California nebula to celebrate our come back on Earth! It was a pleasure working with you, thanks a lot for this opportunity!

Problems Encountered: N/A


Solar Features Observed: N/A

Images submitted with this report: N/A

Problems Encountered: N/A


Sol Summary – March 13th

Crew 223 Sol Summary Report 13-03-2020

Sol: 12 Summary Title: That’s all folk’s

Author’s name: Aurélien Mure – Crew Commander

Mission Status: crew out of sim

Sol Activity Summary:

The crew woke up at 7:50am. The sport session began at 8:00am. Members are struggling to sleep, that is why we moved the sport session, knowing that the weather did not allow us to go outside in the morning. We did all the cleaning work on the hab, the GreenHab, the science dome and the RAM. As we had to Unmount our LOAC and MegaAres, the crew worked on the Human Factor experiment in the late morning.

We went out with a lot of wind and after 4 weeks of hard labour, the weather station fell down one hour before we were supposed to unmount it. but no worries, nothing is broken. The EVA was not so easy because of the wind.

We washed all the materials and we went out of sim. After a long debriefing on how we have lived the mission, we were ready to cook for tonight !

Look Ahead Plan : Unmount the outdoor experiments.

Anomalies in work: None.

Weather: rainy in the morning, 4°C at 8:30am then cloudy and windy in the afternoon.

Crew Physical Status: All crew members safe and sound.

EVA: LOAC and MegaAres unmounted

Reports to be filed: Journalist report, Operations report, Green Hab report, EVA report, Astronomy report, Scientist, food inventory, Mission summary

Support Requested: None

Operations Report – March 13th

Crew 223 Operations Report 13/03/2020

SOL: 12

Name of person filing report: Luc FORTIN

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Thursday generator:

Off at 6:30am

On at 10:16pm SOC was 89%

Friday generator:

Off at 6:00am

On at 10:00pm SOC was XX%

Diesel Reading – 60%

Station Propane Reading – 70%

Ethanol Free Gasoline – 5.5 gallons

Water (loft tank): 47 gallons

Water Meter: 149826.48 units

Water (static tank) – 202 gallons

Static to Loft Pump used – yes

Water in GreenHab – 110.3 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Perseverance rover used: yes

Hours: 152.4

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Nothing to report

Beginning charge: Nothing to report

Ending charge: Nothing to report

Currently charging: Nothing to report

Spirit rover: yes

Hours: 134.2

Beginning charge: 100%

Ending charge: 92%

Currently charging: yes

Opportunity rover used: no

Hours: 70.3

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Curiosity rover used: no

Hours: 139.6

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: yes (1, 2 and 3)

Reason for use: ATVs 1, 2, 3 started during the EVA Engineer.

Oil Added? no

ATV Fuel Used: Nothing to report

Hours the ATVs were Used today: Nothing to report

Notes on ATVs: Nothing to report.

HabCar used and why, where? The HabCar is parked at the MDRS Campus.

CrewCar used and why, where? The CrewCar is in Grand Junction.

General notes and comments: Nothing to report

Summary of internet: All ok.

Summary of suits and radios: Suits 7, 8, 9, 10 were used during the EVA.

Summary of Hab operations: Nothing to report.

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

EVA Report – March 13th

Crew 223 EVA Report 13-03-2020

EVA # 12

Author: Clément Plagne (Journalist)

Purpose of EVA: Dismantling experiments LOAC and MegaARES

Start time: 15:05

End time: 15:43

Narrative: The experiments were dismantled quickly and efficiently in spite of the heavy winds and wet ground.

Destination: LOAC and MegaARES area, off Entrance Road

Coordinates (use UTM NAD27 CONUS): 12S 4250800 518500

Participants: Clément Plagne (Journalist, EVA Leader), Aurélien Mure (Commander), Luc Fortin (Engineer), Valentin Bellemin-Laponnaz (GreenHab Officer)

Road(s) and routes per MDRS Map: Entrance Road, Cow Dung Road

Mode of travel: Rovers : Spirit, Perseverance

GreenHab Report – March 13th

Crew 223 GreenHab Report – 13-03-2020 GreenHab Officer: Valentin BELLEMIN-LAPONNAZ

Environmental control: Ambient with Door open

Average temperature: 24°C

Low temperature: 23°C
High temperature: 25°C

Hours of supplemental light: 4 hours (from 9P.M to 1 a.m.)

Daily water usage for crops: 12.7 gallons

Daily water usage for research and/or other purposes: 0 gallon

Water in Blue Tank— 97.5 gallons

Time(s) of watering for crops: 3 (one only for tomatoes)

Changes to crops: NA

Narrative: This morning, was a special as the weather was windy and rainy. I watered the GreenHab as usual. I felt a bit sad as it is the end of the mission and many fruits will be to be good to eat in two weeks (approx. 15 cucumbers, 5-6 tomatoes). I am also sad to learn that no crew is coming next week. Yet, I had a really good times in the GreenHab. I enjoyed a lot taking care of the plants. I was like on heaven.

Harvest: 25g Basil / 5g Oregano / 20g Sage / 100g Rocket

Support/supplies needed: NA

Journalist Report – March 13th

Crew 223 Journalist Report 13Mar2020

Author: Clément Plagne, Journalist

Sol 12: Closing in

It’s weird to wake up knowing that that heavy airlock door you heard shut such a short time ago was going to open just that evening. Honestly, while knowing that it’s over will be a relief, I think all of us would also like it to last just a little longer. It’s part of the experience to know that you can never do everything you intend to, and that you will face challenges that will slow you down. The job is as much about dealing with these challenges as it is doing what you came here to do. And challenges we had to face, even today. We had so many great days early on, I think that the MDRS gods decided to punish us before the end.

We awoke not only to the thought of that door, but also to pouring rain. A bad situation on clean-up day. The corridors between buildings in the station are insulated, but we still walk on the ground, or, as we had to today, the mud. Wind was howling even after rain had stopped, and the EVA that we had already moved to the afternoon was becoming even more perilous. After four weeks of loyal services to our two crews, the weather station for the LOAC experiment finally lost against the wind and fell over, mere hours before being dismantled. Nevertheless, the EVA was a great demonstration of all we’d learned during our previous expeditions. In hostile conditions of strong winds and loose, muddy terrain, we managed to perform all intended procedures quickly, safely and efficiently. The conclusion of a job well done on all EVAs.

And, some time later, the experiments were all packed up and ready to go back home. We’d been entrusted with them, and we hope we did a great job of gathering data. Everyone in the station did their best, that is, excellent work, so we have high hopes for the results to be valuable. We were just counting minutes until we could say the simulation was broken, and there it was: the outside, not seen through a glass helmet. We all ran around in the mud like children, and waited for Shannon, whom we’d invited over for dinner, to come inside the Hab. She’s done a fantastic job of helping us throughout the mission, and we are all grateful to have her here.

I’m writing these last few lines while she’s here: she has far too many good stories about her tenure at MDRS, and I’m finding myself hurriedly doing my best to finish up before missing the comms window. Tomorrow will be my last report, and there’s a lot to think about and a good look back to do.

Copyright © The Mars Society. All rights reserved. | Main Site